A ladder is resting against a wall.The top of the ladder touches the wall at a height of 9 feet.Find the length of the ladder if the length is 3 feet more than its distance from the wall.
Answer by mananth(16949)

A ladder is resting against a wall.The top of the ladder touches the wall at a height of 9 feet.Find the length of the ladder if the length is 3 feet more than its distance from the wall.
Let the distance of the ladder from the wall be x
The length of the ladder = x+3
The ladder touches the wall at a vertical height of 9 feet.
(x+3)^2 = x^2 + 9^2 ( using the pythagoras theorem
x^2 + 6x+ 9 = x^2 + 81
x^2- x^2 +6x = 81-9
6x= 72
x= 12
Therefore the length of the ladder will be x+3 = 12 + 3 =15 Answer.
M Ananth----------- mananth@hotmail.com